OpenStudy (anonymous):

HELP ASAP. a social network site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The information represents this information, where X represents the total number of people who became members of the site that month x=0.8x+72 how many people who became members of the site were between 30 and 40 year of age? 112 288 360 432

OpenStudy (anonymous):

0.2x=72 x=72/02.=360 360(0.8)=288 people 30-40yrs old

OpenStudy (anonymous):

ok thanks ill let u know if i need help with any more

OpenStudy (anonymous):

but r u sure

OpenStudy (anonymous):

can u please give me madal?

OpenStudy (anonymous):

medal

OpenStudy (anonymous):

how do u do it

OpenStudy (anonymous):

see were my answer is?

OpenStudy (anonymous):

good thank you

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i got one or two more questions can u help with them

OpenStudy (anonymous):

post it on questions and I will meet u there

OpenStudy (anonymous):

solve for X 4/5x+4/3=2x

OpenStudy (anonymous):

if u want I can do it there?

OpenStudy (anonymous):

yes plz